can you help me to solve for x in terms of y and solve for y in terms of x for the following equation.
3y^2+4xy-9y^2=-1
Email - acrown1@umbc.edu

Answer by stanbon(75887)

can you help me to solve for x in terms of y and solve for y in terms of x for the following equation.
3y^2+4xy-9y^2=-1
Combining like terms you get the following:
4xy-6y^2=-1
Then x=[6y^2-1]/4y
To solve for "y" rearrange and use the quadratic formula to get:
6y^2-4xy-1=0
y=[4x+sqrt((4x)^2+4(6)]/12=[4x+sqrt(16x^2+24)]/12=[2x+sqrt(x^2+6)]/6
OR
y=[2x-sqrt(x^2+6)]/6
Cheers,
Stan H.